Number of moles of sodium carbonate in 250cm3


1
Expert's answer
2021-03-04T06:38:49-0500

n = m/M

M (Na2CO3) = 106 g/mol

density = m/V

m = density x V

density (Na2CO3) = 2.54 g/cm3

m (Na2CO3) = 2.54 x 250 = 635 g

n (Na2CO3) = 635 / 106 = 6 mol